what are the types of transformations? how do you determine them?
ex) y = 1/2 x - 4 ---> y = -1/2 x -4

Answer by stanbon(75887)

what are the types of transformations? how do you determine them?
ex) y = 1/2 x - 4 ---> y = -1/2 x -4
------
Notice that the change from (1/2)x to (-1/2)x only
changes the slope of the line. Since the -4
term has not changed, the y-intercept has not changed.
---
Your transformation could be written this way
y = (1/2)x-4 to y = -((1/2)x-4)-8
That would described as 2 transformations.
1st::: A reflection in the x-axis.
followed by
2nd: A vertical shift down of -8.
===============
cheers,
Stan H.